39.8g of a mixture of potassium chloride,kcl, and potassium trioxochlorate were heated to a constant mass. If the residue weighed 28.9g .what was the percentage mass of the potassium chliride in the mixture? (K39, CL=35.5,O=16) hint KCL is not decomposed on heating. Pls in need the answer
1
Expert's answer
2019-10-02T08:03:49-0400

Let the mass of "KClO_3" be x gram

Reaction of decomposition is given by:

"2KClO_3\to 2KCl+ 3O_2"

2 moles of Potassium trioxochlorate produces 2 moles of Potassium chloride

So,

"\frac{x}{122.5}" Moles will produce same amount of KCl

"\\""\frac{x}{122.5}\ mole \ KCl=\frac{x}{122.5}\times 74.5\ g\ KCl"

Now final mass =initial mass of KCl +mass of KCl produced due to decomposition = "28.9\ g"

So,

"39.8-x\ +\frac{74.5x}{122.5}=28.9"

"0.391836x=10.9\\x=27.8177\ g"

Initial masso f Potassium chloride = "38.9-x=11.0823 \ g\ KCl"

Percentage composition of KCl = "\frac{11.0823}{39.8}\times100=27.844\%"